# masers — ISM: individual objects (G33.641-0.228) —stars: formation — stars: massive

### v1

@@ 3.5 Properties of the flux and circular polarization variability
  • The circular polarization does not match the flux density variability. When the large burst occurred in 2014 and the flux density increased, the polarized flux density also increased, but the polarization decreased to 0. On the other hand, 2016 intensive monitoring showed no systematic correlation of circular polarization and flux density.
@@ 4 Discussions
In the data reduction, we have assumed that the circular polarization of component III is zero. The derived circular polarization of components I and IV based on this assumption was zero within the error. This does not prove, but it supports the assumption that the circular polarization of component III is zero.
@@
Circularly polarized radiation has been detected only in component II, and the bursts also occur only in this component. Therefore, the burst and the circularly polarized radiation could be related to, and/or caused by a common physical mechanism. Linear polarization often arises from interstellar matter along the propagation path. If dust and magnetic fields exist along the propagation path of visible light from stars, linear polarization is generated by the alignment of the dust (e.g., Heiles 2000). This mechanism, however, is unlikely to produce circular polarization during the propagation of 6.7 GHz radio emission.
When maser emission occurs in a region permeated by a magnetic field, differences in the intensities of the circularly polarized components arise, and at the same time, the Zeeman effect produces a frequency (or velocity) shift between the two circular polarizations. Vlemmings (2008) observed methanol masers in G33.641-0.228 and found a line-of-sight velocity difference of $\Delta v=-0.89\pm 0.17$ m s-1. We performed Gaussian fitting to the RHCP and LHCP line profiles separately to measure their peak velocities. Using all data obtained between 2009 and 2015, we calculated the velocity differences between the peak velocities of the two circular polarizations and derived the mean values and standard errors. As a result, the velocity difference for component II is $\Delta v=-1.1\pm 0.5$ m s-1, which is consistent, within the uncertainties, with the value obtained by Vlemmings (2008). The magnetic field strength inferred from this velocity difference is $22\pm 10$ mG when adopting the conversion coefficient of 49-49 m s-1 G-1 used by Vlemmings (2008), or $260\pm 120$ mG when using the coefficient of 4.6-4.6 m s-1 G-1 derived by Takagi et al. (2021). The Zeeman coefficients of the 6.7 GHz methanol maser have also been derived by Lankhaar et al. (2018) for different hyperfine components. If the observed line-of-sight velocity difference is caused by hyperfine structure, the magnetic field strengths derived above should be regarded as upper limits to the true values. The derived magnetic field strength is comparable to those of methanol masers in other sources (Vlemmings 2008), and it is therefore unlikely that the unusually large circular polarization observed in this source arises solely from an exceptionally strong magnetic field acting only on this particular component.
@@
When circularly polarized emission occurs behind a maser cloud, the amplified maser emission also becomes circularly polarized. The circular polarization observed in G33.641-0.228 may have been amplified by the circularly polarized continuum radiation behind it. There are two well-known mechanisms for circularly polarized continuum radiation in solar radio emission. One is the gyro-synchrotron radiation observed in solar radio bursts and radio flares of T Tauri stars (Dulk 1985; Phillips et al. 1993), where high-energy charged particles interact with the magnetic field and emit circularly polarized radiation. The other is gyro-resonance radiation, in which thermal plasma within a strong magnetic field emits circularly polarized radiation (Shibasaki et al. 1994). In YSOs, circularly polarized radiation has been observed during flares of T Tauri stars in the microwave band (White et al. 1992; Phillips et al. 1993; Skinner & Brown 1994; Phillips et al. 1996). It is known that magnetic fields are involved in burst-like phenomena in YSOs, e.g., Takasao et al. (2019) conducted three-dimensional magnetohydrodynamic simulations to explain X-ray flares in YSOs and showed that the flares are caused by magnetic reconnection.
@@
We adopt a working hypothesis that an explosive event similar to a solar radio burst occurred on the surface of the protostar and the circularly polarized continuum radiation was generated, then the radiation was amplified by the maser cloud of component II along the line of sight from the explosive event region to the observer. In this hypothesis, the flux variability of the maser burst is not due to the variability of the physical property of the maser cloud, but due to the variability of the background radiation. The explosive event region of the protostar, the component II maser cloud, and the observer are aligned linearly, while other maser clouds are not. Bartkiewicz et al. (2011, 2012) investigated the spatial distributions of the methanol and water masers in G33.641-0.228. The methanol masers exhibit an elliptical or arc-like distribution, while the water masers are distributed in two opposite directions roughly perpendicular to the methanol maser structure. The methanol masers are thought to primarily trace an accretion disk, whereas the water masers trace an outflow driven by a protostar. It is therefore inferred that the protostar is located at the position where the distributions of the methanol and water masers intersect. This position is close to the western edge of the methanol maser distribution and is nearly coincident with the location of the bursting maser spots of component II (Bartkiewicz et al. 2011, 2012; Fujisawa et al. 2012; note that the methanol maser was not in a bursting phase at the time of the VLBI observations). Consequently, it is possible that the protostar is located behind the bursting maser spots.
Based on this working hypothesis, we attempt to explain the observed properties of the burst and the circular polarization in the following.
@@
  • Property 1: Short-term flux variation. The rise and fall of the flux, repeating with short intervals, are observed in the solar radio bursts (microwave IV bursts). The timescale of solar radio bursts is 0.01 d (Dulk 1985). The time scale of the flux variability of the maser bursts is 30 times larger than that of solar radio bursts, but other qualitative properties are very similar. Note that radio bursts with circular polarization have also been observed in the low-mass pre-main-sequence stars T Tauri and V773 Tau, with variability timescales of less than a few days for T Tau and a few hours for V773 Tau (White et al. 1992; Phillips et al. 1993; Skinner & Brown 1994; Phillips et al. 1996).
@@
  • Property 2: Circular polarization and its variability. In solar radio bursts, circular polarization due to either gyro-synchrotron radiation or gyro-resonance radiation is observed and fluctuates over short periods of time. The circular polarization degree is about 10% for gyro-synchrotron radiation (Dulk 1985). If a similar event occurred in G33.641-0.228, the observed fluctuations in circular polarization degree can be explained.
@@
  • Property 3: The correlation between flux density and polarization changes from day to day (e.g. Figure 9) for the maser burst. During the large burst, circularly polarized flux increased, but polarization decreased (Figure 8). The observed decrease in the degree of circular polarization with increasing flux density (anti-correlation) may be explained as follows. The radiative transfer equation for maser radiation is
    $dI_{\rm{R/L}}/ds=\kappa(I)I_{\rm{R/L}},$
    where $I_{\rm{R/L}}$ are the intensity of the maser at RHCP and LHCP, and the maser opacity $\kappa(I)$ is a function of the total intensity ($I=I_{\rm{R}}+I_{\rm{L}}$). In the unsaturated regime, $\kappa(I)=\kappa_{0}$, the polarized components are
    $I_{\rm{R/L}}=\exp(\tau)B_{\rm{R/L}},$
    where $\tau$ is the optical depth of the maser and $B_{\rm{R/L}}$ are the brightness of the incident background radiation at RHCP and LHCP. Computing the polarization fraction
    $\Pi_{\rm{c}}=(I_{\rm{R}}-I_{\rm{L}})/(I_{\rm{R}}+I_{\rm{L}})=(B_{\rm{R}}-B_{\rm{L}})/(B_{\rm{R}}+B_{\rm{L}})=\Pi_{\rm{c}}^{\rm{B}}$
    yields the background field polarization fraction ($\Pi_{\rm{c}}^{\rm{B}}$). If we instead take the case of a saturated maser (assuming high saturation, $I\gg I_{\rm{sat}}$, where $I_{\rm{sat}}$ is the saturation intensity, and using that $B_{\rm{R}}$, $B_{\rm{L}}\ll I_{\rm{sat}}$), we let $\kappa(I)=\kappa_{0}/[1+I/I_{\rm{sat}}]$, the polarization fraction is
    $\Pi_{\rm{c}}\sim\Pi_{\rm{c}}^{\rm{B}}/(I/I_{\rm{sat}}),$
    where $I/I_{\rm{sat}}$ is the saturation degree. This function is inversely correlated with the intensity. This could explain the observation described in 3.2.2, and the anticorrelation between the flux density and polarization fraction.
@@
  • Property 6: On October 9, 2012, a strong negative circular polarization of 0.293-0.293 was observed. This requires an ad hoc explanation: that day, a reversal of the circular polarization coincidentally occurred on the surface of the protostar.
Taken together, the working hypothesis can explain the generation of circular polarization and short-term flux variability. The anti-correlation between flux density and polarization could be explained by the saturation model. However, the presence of strong circular polarization even during the quiescent period, the stability of positive circular polarization over 7 years, and the occurrence of a large negative circular polarization observed only once remain unexplained.
@@
If VLBI observations are made during bursts, and the location where the bursts occur will be identified, leading to an understanding of the burst mechanism. Additionally, if the background continuum emission is amplified by masers, the detection of this continuum emission would be a clue. It is expected that the continuum emission is circularly polarized and shows the same light curve with the maser, and that the continuum emission region exists only behind the component II maser cloud and not behind the other maser clouds. The size of the continuum emission region could be estimated from the 0.3 d variability timescale, which corresponds to a size smaller than 50 au or 13 mas and can be easily measured by VLBI. If such continuum emission is detected, the burst mechanism can be clarified, consequently the amplification rate of the maser can be directly measured.
Fujisawa et al. (2012) and Bartkiewicz et al. (2009, 2012) reported VLBI observations of the 6.7 GHz methanol maser in G33.641-0.228, but these observations were not carried out during a bursting period. Unfortunately, no burst has been observed since 2022 by the long-term methanol maser monitoring program of Hitachi 32-m telescope [^2] Ibaraki 6.7 GHz class II methanol maser database (iMet) as shown in Figure 11.
@@ 5 Conclusion
The 6.7 GHz methanol maser in the high-mass star-forming region G33.641-0.228 shows burst-like flux variability. We have carried out monitoring observations to investigate the mechanism of this burst with the Yamaguchi 32-m radio telescope from 2009 to 2016. The data were analyzed independently for both left and right circular polarization to examine the circular polarization properties of the burst.
Through these observations, the following facts have been revealed. After the large burst, a short-term rise and fall of flux density with a time scale (e-folding time) of 0.3 d was observed. Among the spectral components, only component II, which showed the burst, exhibited large circular polarization. Circular polarization was observed from 2009 to 2016, and it ranged roughly from 0.0 to 0.2 and varied on a time scale of less than one day. When a burst occurred, the circularly polarized flux increased with the averaged flux density, but the circular polarization tended to decrease when the flux density increased. The variability of the flux density and the circular polarization did not show any clear correlation during the low-flux-density period. Only once a large negative circular polarization of -0.293 was observed during the observation period.
We proposed a model to explain these properties, in which an explosive event similar to a solar radio burst occured on the surface of the protostar and the generated radiation with high circular polarization is amplified by the component II maser cloud on the line of sight from the location of the explosive event to the observer. This model can explain the generation of large circular polarization and short-term flux variations. Khaibrakhmanov et al. (2025) proposed a physical model involving reconnection between the protostellar magnetosphere and the disk magnetic field to explain the observed short-term flux variability.
@@
This work was partially supported by the Inter-university collaborative project, the Japanese VLBI Network (JVN) of the National Astronomical Observatory of Japan. The authors thank Dr. Hachisuka, Dr. Matsumoto, Mr. Takase, Ms. Fukui, and Ms. Koga for their assistance with the observations. KF thanks Dr. Kazi Rygl and Dr. Yuanwei Wu for providing the distance information. The authors thank Dr. Ross A. Burns for proofreading the manuscript. The authors also thank the anonymous referee for many helpful comments. The possible explanation for the anti-correlation between flux density and polarization was suggested by the referee. The authors also thank KDDI Corporation for their support of the Yamaguchi 32-m radio telescope.

### v2

@@ 3.5 Properties of the flux and circular polarization variability
  • The circular polarization does not match the flux density variability. When the large burst occured in 2014 and the flux density increased, the polarized flux density also increased, but the polarization decreased to 0. On the other hand, 2016 intensive monitoring showed no systematic correlation of circular polarization and flux density.
@@ 4 Discussions
In the data reduction, we have assumed that the circular polarization of component III is zero. The derived circular polarization of components I and IV based on this assumption was zero within the error. This does not prove but supports the assumption that the circular polarization of component III is zero.
@@
Circularly polarized radiation has been detected only in component II, and the bursts also occur only in this component. Therefore, the burst and the circularly polarized radiation could be related and/or caused by a common physical mechanism. Linear polarization often arises from interstellar matter along the propagation path. If dust and magnetic fields exist along the propagation path of visible light from stars, linear polarization is generated by the alignment of the dust (e.g., Heiles 2000). This mechanism, however, is unlikely to produces circular polarization during the propagation of 6.7 GHz radio emission.
[…]
It is well established that magnetic fields are involved in burst-like phenomena in YSOs. Takasao et al. (2019) conducted three-dimensional magnetohydrodynamic simulations to explain X-ray flares in YSOs and showed that the flares are caused by magnetic reconnection. Although X-ray flares differ from the maser bursts in G33.641-0.228, magnetic fields may likewise play a role in the maser bursts.
@@
When circularly polarized emission occurs behind a maser cloud, the amplified maser emission also become circularly polarized. The circular polarization observed in G33.641-0.228 may have been amplified by the circularly polarized continuum radiation behind it. There are two well-known mechanisms for circularly polarized continuum radiation. One is gyro-synchrotron radiation observed in solar radio bursts (Dulk 1985), where high-energy charged particles interact with the magnetic field and emit circularly polarized radiation. The other is gyro-resonance radiation, in which thermal plasma within a strong magnetic field emits circularly polarized radiation (Shibasaki et al. 1994). In YSOs, circularly polarized radiation has been observed during flares of T Tauri stars in the microwave band (Skinner & Brown 1994; Phillips et al. 1996).
@@
We adopt a working hypothesis that an explosive event similar to solar radio burst occurred on the surface of the protostar and the circularly polarized continuum radiation was generated, then the radiation was amplified by the maser cloud of component II on the line of sight from the explosive event region to the observer. In this hypothesis, the flux variability of the maser burst is not due to the variability of the physical property of the maser cloud, but due to the variability of the background radiation. The explosive event region of the protostar, the component II maser cloud, and the observer are aligned linearly, while other maser clouds are not. Based on this working hypothesis, we attempt to explain the observed properties of the burst and the circular polarization in the following.
@@
  • Property 1: Short-term flux variation. The rise and fall of the flux, repeating with short intervals, are observed in the solar radio bursts (microwave IV bursts). The timescale of solar radio bursts is 0.01 d (Dulk 1985). The time scale of the flux variability of the maser bursts is 30 times larger than that of solar radio bursts, but other qualitative properties are very similar. Note that radio bursts with circular polarization have also been observed in the low-mass pre-main-sequence stars T Tauri and V773 Tau, with variability timescales of less than a few days for T Tau and a few hours for V773 Tau (Skinner & Brown 1994; Phillips et al. 1996).
@@
  • Property 2: Circular polarization and its variability. In solar radio burst, circular polarization due to either gyro-synchrotron radiation or gyro-resonance radiation is observed and fluctuates over short periods of time. The circular polarization degree is about 10% for gyro-synchrotron radiation (Dulk 1985). If a similar event occurred in G33.641-0.228, the observed fluctuations in circular polarization degree can be explained.
@@
  • Property 3: The correlation between flux density and polarization changes from day to day (figure 9) for the maser burst. During the large burst, circularly polarized flux increased, but polarization decreased (figure 8). There is no simple explanation for these properties.
@@
  • Property 6: On October 9, 2012, a strong negative polarization of 0.293-0.293 was observed. This requires an ad hoc explanation that, on that day, a reversed circular polarization coincidentally occurred on the surface of the protostar.
Taken together, the working hypothesis can explain the generation of circular polarization and short-term flux variability, while it does not well explain the correlation between flux density and polarization, the presence of strong circular polarization radiation even in the quiet period, the stability of positive circular polarization for 7 years, and the occurrence of a large negative circular polarization observed only once.
@@
If VLBI observations will be made during bursts, and the location where the bursts occur will be identified, leading to an understanding of the burst mechanism. Additionally, if the background continuum emission is amplified by masers, the detection of this continuum emission would be a clue. It is expected that the continuum emission is circularly polarized and shows the same light curve with the maser, and that the continuum emission region exists only behind the component II maser cloud and not behind the other maser clouds. The size of the continuum emission region could be estimated from the timescale of the variability of 0.3 d, which would be smaller than 50 au or 13 mas and can be easily measured by VLBI. If such continuum emission will be detected, the burst mechanism can be clarified, consequently the amplification rate of the maser can be directly measured.
Fujisawa et al. (2012) and Bartkiewicz et al. (2009, 2012) reported VLBI observations of the 6.7 GHz methanol maser in G33.641-0.228, but these observations were not carried out during the bursting period. Unfortunately, no burst was observed after 2022 by the long term methanol maser monitoring program of Hitachi 32-m telescope [^2] Ibaraki 6.7 GHz class II methanol maser database (iMet) as shown in figure 11.
@@ 5 Conclusion
The 6.7 GHz methanol maser in the high-mass star-forming region G33.641-0.228 shows burst-like flux variability. We have made monitoring observations to investigate the mechanism of this burst with the Yamaguchi 32-m radio telescope from 2009 to 2016. The data were analyzed independently for both left and right circular polarization to examine the circular polarization property of the burst.
Through this observation, the following facts have been revealed. After the large burst, a short-term rise and fall of flux density with a time scale (e-folding time) of 0.3 d were observed. Among the spectral components, only component II, which showed the burst, exhibited large circular polarization. Circular polarization was observed from 2009 to 2016, and it ranged roughly from 0.0 to 0.2 and varied on a time scale of less than one day. When a burst occurred, the circularly polarized flux increased as the averaged flux density, but the circular polarization tended to decrease when the flux density increased. The variability of the flux density and the circular polarization did not show any clear correlation at the low flux density period. Only once a large negative circular polarization of -0.293 was observed during the observation period.
We proposed a model to explain these properties, in which an explosive event similar to a solar radio burst occurs on the surface of the protostar and the generated radiation with high circular polarization is amplified by the component II maser cloud on the line of sight from the location of the explosive event to the observer. This model can explain the generation of large circular polarization and short-term flux variations. Khaibrakhmanov et al. (2025) proposed a physical model of reconnection between the protostellar magnetosphere and the disk magnetic field to explain the observed short-term flux variability.
@@
This work is partially supported by the Inter-university collaborative project, the Japanese VLBI Network (JVN) of the National Astronomical Observatory of Japan. The authors thank to Dr. Hachisuka, Dr. Matsumoto, Mr. Takase, Ms. Fukui, and Ms. Koga for their assistance with the observations. KF thanks to Dr. Kazi Rygl and Dr. Yuanwei Wu for providing the distance information. The authors also thank to KDDI Corporation for their support of the Yamaguchi 32-m radio telescope.